# Bistability of electron temperature in atomically thin semiconductors in the presence of exciton photogeneration

### 07 Mar 2026 v1

@@ I Introduction
In this paper, we investigate nonlinear effects [26, 27, 28, 29, 30, 31, 32] arising from the conversion of excitons to trions in TMDC MLs under continuous exciton photogeneration. We show that the exponential increase in trion dissociation rate with temperature leads to a stepwise change in the ratio of the density of unbound resident charge carriers to the trion density. Combined with the fact that trions (being heavier than free carriers) exhibit lower conductivity and shorter energy relaxation times [33], this effect gives rise to bistability [34] in the system’s steady state when it is heated by an external low-frequency field. The first state is a low-temperature one, in which most of the resident charge carriers are bound to excitons, forming trions. The second is a high-temperature state, in which most trions are dissociated, leading to more efficient heating. Switching between these two states occurs on a timescale of 10–100 ps and is accompanied by abrupt changes in various system observables—such as temperature, electric current, fraction of the absorbed-transmitted external electromagnetic field and the intensity of exciton or trion luminescence—forming a hysteresis loop as the heating power is varied.
@@ II Dynamic equations of the system
[…]

### 24 Mar 2026 v2

@@ I Introduction
In this paper, we investigate nonlinear effects [26, 27, 28, 29, 30, 31, 32] arising from the conversion of excitons to trions in TMDC MLs under continuous exciton photogeneration. Namely, the phenomenon of bistability [33], which occurs in a wide range of systems [34, 35, 36, 37, 38, 39, 40, 41], from classical oscillators to metasurfaces [42]. We show that the exponential increase in trion dissociation rate with temperature leads to a stepwise change in the ratio of the density of unbound resident charge carriers to the trion density. Combined with the fact that trions (being heavier than free carriers) exhibit lower conductivity and shorter energy relaxation times [43], this effect gives rise to bistability in the system’s steady state when it is heated by an external low-frequency field. The first state is a low-temperature one, in which most of the resident charge carriers are bound to excitons, forming trions. The second is a high-temperature state, in which most trions are dissociated, leading to more efficient heating. Switching between these two states occurs on a timescale of 10–100 ps and is accompanied by abrupt changes in various system observables—such as temperature, electric current, fraction of the absorbed-transmitted external electromagnetic field and the intensity of exciton or trion luminescence—forming a hysteresis loop as the heating power is varied.
@@ II Dynamic equations of the system
Figure 1: Diagram of the processes described by Eq. (1): (a) Impact trion dissociation with the rate $\beta n_{T}n_{e}$ caused by electrons with an energy on the order of $E_{T}$; (b) Radiative recombination of an exciton within a trion ($n_{T}/\tau_{T}$), during which an optical photon is emitted and the outor electron remains in the conduction band; (c) Exciton energy relaxation via electron capture from the conduction band, accompanied by optical phonon emission ($\gamma n_{e}n_{X}$); (d) Exciton photogeneration $G_{X}$ and radiative recombination ($n_{X}/\tau_{X}$), see text for details. (f) The dependence of the density of unbound electrons $n_{e}$ and trion density $n_{T}$ on the temperature of the electron gas. The solid lines correspond to the solution of Eq. (4), while the dashed lines correspond to the Saha equation Eq. (5). Inset shows dependence of ratio $n_{T}/n_{e}$ on the temperature of the electron gas. The parameters of calculation are $G_{X}=4.5\cdot 10^{10}$ cm-2ps-1, $\gamma=0.1$ cm2s-1, $n_{c}=10^{12}$ cm-2, $\tau_{X}=\tau_{T}=50$ ps, $\beta=w_{0}\frac{k_{B}T}{E_{T}}e^{-\frac{E_{T}}{k_{B}T}}$, with $w_{0}=1.16\cdot 10^{3}$ cm2s-1, $E_{T}=25\penalty 10000\ \text{meV}$ [19, 23, 24].